# Emergent Turbulence in Nonlinear Gravity

### v1

@@ Abstract
Gravity in nonlinear and dynamical regimes underpins spectacular astrophysical phenomena and observable consequences, from the early universe to black hole collisions. In these extreme environments, inverse energy cascades — mediated by nonlinear interactions — may help explain the near scale-invariance of cosmic structure and the simplicity of gravitational waves from binary black hole mergers. Yet the presence, characteristics, and generality of such interactions in full General Relativity remain largely unexplored. Here we show that two types of nonlinear interactions — a four-mode and a three-mode interaction — emerge in the fully nonlinear regime, and can indeed channel inverse energy cascades by inducing resonant and anti-damping instabilities. This establishes what was previously only hinted at in highly specialized perturbative contexts. We further demonstrate a “laminar” to “turbulent” transition for the largest-possible angular structure in General Relativity, whereas finer structures remain persistently turbulent. Our results reveal the impact and generality of these nonlinear interactions (instabilities), which can be key to understanding observations ranging from cosmological to kilometer scales. We anticipate that our work will shed new light on nonlinear gravitational phenomena and their consequences, such as constructing gravitational wave templates and testing General Relativity in the most extreme regime. Moreover, our work is a starting point for addressing nonlinear gravitational interactions using ideas and methods inspired by fluid dynamics.
@@ introduction
Understanding truly nonlinear phenomena presents significant challenges. One particular example is furnished by fluid dynamics. Its significant degree of complexity requires ingenious approaches to simplify the task of unraveling possible phenomenology. In this context, a powerful approach has been to consider systems that are continuously driven by an external source. Such driving allows for the establishment of a steady-state regime in which systematic insights can be gained. In particular, it enables the identification of a direct energy cascade in 3+1 dimensions and an inverse cascade in 2+1 dimensions [1, 2, 3]. This understanding has provided new avenues to both control and capture the complex dynamics exhibited by turbulent regimes, e.g. [4, 5, 6, 7].
@@
Gravity, as described by General Relativity, remains only partially understood in the fully nonlinear regime. While major results have been obtained for stationary spacetimes, asymptotic structures, and perturbations around special solutions, the nonlinear and dynamical aspects still pose significant challenges — even though numerical simulations can now routinely model the collisions of compact binary systems and predict the resulting gravitational-wave signatures.
@@
Our main interest here is to unravel this nonlinear regime, particularly in 3+1 dimensions. Hints of exciting phenomena have been obtained through perturbative analyses in convenient settings where linear perturbations persist long enough to channel interesting nonlinear behavior. Notably, [8, 9] uncovered a parametric instability that preferentially transfers energy to larger scales. These works, in turn, were motivated by the “fluid-gravity duality”, in which certain gravitational systems in d+1 dimensions (within asymptotically anti-de Sitter spacetimes) are dual to (d-1)+1 relativistic hydrodynamics [10]. In this framework, long-wavelength perturbations of black holes correspond, on the hydrodynamic side, to scenarios where turbulence can take place [11, 12]. Explicit illustrations of such behavior have been produced through numerical simulations on both sides of the duality [13, 14]. However, such duality requires the unphysical assumption of gravity with a negative cosmological constant.
These observations signal that a new chapter in the exploration of nonlinear gravity is ready to unfold. Here, important lessons can be drawn from fluid dynamics and its turbulent phenomena. There is, however, a crucial difference: in fluid dynamics, nonlinear interactions are driven by the principal part of the equation — in particular, the advective terms, which depend on the velocity field itself, may cause physical characteristics to cross. In General Relativity, such crossing does not occur; instead, nonlinear phenomena are governed by lower-order terms. In simple terms, gravity “gravitates” — it can focus energy and mediate interesting behavior. Taking a page from the strategy followed in fluid dynamics, we formulate this exploration by introducing a steady driving source to gravitationally stir spacetimes of interest with incoming gravitational waves, and then capturing the full nonlinear behavior through numerical simulations of Einstein’s equations. These driving waves have prescribable amplitude, frequency, and spatial extent, allowing us to systematically probe the spacetimes’ response. By adjusting the amplitude, we can infer the type of mode interactions underlying the observed behavior; by varying the frequency and angular distribution, we can trace the direction of energy cascades.
Our analysis enables a clear distinction between linear (laminar) and nonlinear (turbulent) regimes, as well as a classification of transient versus steady-state responses. Moreover, we systematically characterize energy transfer across frequencies and spatial scales, and identify the dominant interaction channels along with their associated instabilities. Together, these insights allow us to construct a comprehensive framework that captures the main mechanisms in fully nonlinear General Relativity that drive inverse cascades and mediate energy transfer toward longer wavelengths.
@@
We “stir” spacetimes by injecting (quasi-)steady gravitational waves through the boundary of a finite simulation domain, as described in Methods. Such driving waves are characterized by their tunable amplitude, angular structure, and frequency — a freedom that is key to unraveling nonlinear mode interactions as well as exploring energy cascades in both temporal and spatial directions.
We study scenarios that initially correspond either to flat spacetime or to a non-spinning, asymptotically flat black hole [^1] Linear perturbations in this regime decay faster than in the spinning case, making it a priori more difficult to excite nonlinear behavior., where the injected waves act as a sustained source of dynamical self-interaction and turbulent gravitational phenomena. The driving is specified through a curvature scalar at the boundary [15, 16, 17, 18]:
@@
In the asymptotically flat black hole case, the black hole is located at the center of the domain with an initial mass $M_{i}$. We adopt
@@
where the injection amplitude is linearly ramped up to $A_{i}$ over a duration $t_{0}=600M_{i}$. This gradual turn-on is essential for stably simulating scenarios with large $A_{i}$. We typically place the outer boundary at $R_{\rm Bdry}=100M_{i}$ and have verified that its location does not affect the overall conclusions. Throughout, the frequency of the injected wave is fixed at a real value $\omega=\kappa/M_{i}$. We find that when $\kappa$ is sufficiently small, $\lesssim{\cal O}(0.3)$, the injected wave cannot pass through the black hole’s potential barrier to be accreted. This behavior, even in the fully nonlinear regime, is largely consistent with predictions from black hole linear perturbation theory (see Supplementary Information for more details). For high frequencies, accretion does take place, and the black hole’s mass increases at a constant rate with a quadratic dependence on $A_{i}$. In the most nonlinear case, the black hole can grow to more than six times its initial mass. For intermediate frequencies, a slower time dependence is observed until the black hole becomes sufficiently massive for the accretion rate to reach its asymptotic, constant behavior.
(a) $\ell=2$ driver
(b) $\ell=6$ driver
Figure 1: Stirring a non-spinning black hole with gravitational-wave drivers. The driving frequency is fixed at $\omega=0.5/M_{i}$. a. Response in $\ell=6$ of $\mathcal{E}$, to a strong $\ell=2$ driver with injection amplitude $A_{i}=1.2\times 10^{-3}$. The response contains three frequency components, $3\omega,2\omega$, and $\omega$, with amplitude envelopes shown in the top three panels. The $3\omega$ and $\omega$ modes initially undergo laminar excitations via third-order couplings (gray-shaded regions). Subsequently, nonlinear excitations of $2\omega$ and $\omega$ emerge, grow, and saturate (orange-shaded regions). The bottom panel shows a time–frequency representation of all three frequency components; the yellow dashed line highlights an inverse cascade: $3\omega\to 2\omega\to\omega$. b. Response across angular harmonics driven by an $\ell=6$ gravitational wave with $A_{i}=4\times 10^{-4}$, measured continuously in time (in blue). Representative time slices are shown as dot-dashed lines. Each harmonic contains two frequency components $2\omega$ and $\omega$. The $2\omega$ component (top panel) arises from quadratic laminar couplings and maintains a stable angular spectrum over time, whereas the $\omega$ mode continuously redistributes its angular spectrum, yielding an inverse cascade toward lower $\ell$ values.
The spacetime geometry is encoded in the metric tensor $g_{\alpha\rho}$ and it is convenient to monitor the behavior of the driven spacetime through the “tendicity” scalar function $\mathcal{E}$, which quantifies the tidal response on a given worldtube [19, 20, 21, 22, 23] and characterizes the local tidal-stretching experienced by orthogonal observers while crossing the worldtube [19]. This scalar is constructed from the Weyl curvature tensor as:
@@
where $n^{a},s^{a}$ are unit normal vectors orthonormal to the spacetime foliation and the worldtube where tidal effects are measured, respectively. In the initially flat spacetime case, we adopt an inner worldtube located at $R_{\Gamma}$; while in the black hole case, we choose the apparent horizon instead. We examine how tendicity’s harmonic components respond to the injected gravitational-wave driver.
(a) $\ell=2$ driver
(b) $\ell=6$ driver
(c) Four-mode coupling induced by the $\ell=2$ driver
(d) Three-mode coupling induced by the $\ell=6$ driver
Figure 2: Examining emergent instabilities in a stirred black hole. The driving frequency is fixed at $\omega=0.5/M_{i}$. a. Responses to an $\ell=2$ driver with various amplitudes (colors), including the $2\omega$ and $\omega$ components in $\ell=6$, as well as $\omega$ in $\ell=4$. Amplitude envelopes are plotted. These modes first undergo laminar excitations via fourth- and third-order couplings (gray-shaded regions). At low driving amplitudes, the structure persists and extends to later times; at higher amplitudes, new emergent excitations appear. Within an intermediate amplitude range, the modes grow exponentially (gray dashed lines) before saturation (open circles). For $2\omega$ in $\ell=6$, the exponential growth rate scales as $A_{i}^{2}$ and the saturation amplitude as $A_{i}^{4}$. b. Responses to an $\ell=6$ driver with various amplitudes. The $\omega$ mode in $\ell=2$ and 44 always grows linearly (dashed lines) regardless of $A_{i}$. The growth rate scales as $A_{i}^{3}$. c. Feynman-type diagram illustrating the four-mode interaction responsible for the emergent excitations in Fig. 2a. d. Feynman-type diagram illustrating the three-mode interaction that drives the linear growth in Fig. 2b.
@@ Results
One main conclusion drawn from our studies, is the inverse energy cascade in both temporal and spatial directions in the nonlinear regime of dynamical spacetimes. Starting with the case containing a non-spinning, asymptotically flat black hole, Fig. 1a illustrates this inverse cascade while injecting an $\ell=2$ gravitational wave, with frequency $\omega=0.5/M_{i}$ and amplitude $A_{i}=1.2\times 10^{-3}$. The envelope of the response in the $\ell=6$ harmonic of $\mathcal{E}$ is shown, decomposed into its constituent frequency components. The harmonic first undergoes two nonlinear excitations at $\omega$ and $3\omega$ (gray region), arising from third-order couplings: $\omega=\omega+\omega-\omega$ and $3\omega=\omega+\omega+\omega$. Their peak amplitudes have been verified to scale as $A_{i}^{3}$. In the weakly driven regime these two frequency components behave in a laminar manner: their temporal evolution reacts instantaneously to the injected wave, remains phase-locked to it, and does not exchange energy with other modes. These features persist over time, and the same modes are excited regardless of the injection amplitude $A_{i}$. See more details below in Fig. 2a and in Supplementary Information.
@@
Remarkably, for sufficiently strong injections, such as the case $A_{i}=1.2\times 10^{-3}$ shown in Fig. 1a, novel nonlinear excitations emerge at $2\omega$ and $\omega$ sequentially in time, as shown by the orange regions in the figure. These two new nonlinear excitations emerge, grow, and saturate after the laminar modes pass their peaks and the injected gravitational-wave driver reaches its constant amplitude. In particular, the $\omega$ component first undergoes a laminar excitation and then a secondary nonlinear emergence, resulting in a two-peaked structure in its temporal evolution. The bottom panel of Fig. 1a further illustrates this phenomenology and provides insight into the underlying dynamics: the yellow dashed line, which connects the dominant laminar mode $3\omega$ to the subsequently emergent nonlinear modes $2\omega$ and $\omega$, indicates an inverse energy cascade towards smaller frequencies.
@@
Figure 1b considers an $\ell=6$ gravitational-wave driver with $\omega=0.5/M_{i}$ and $A_{i}=4\times 10^{-4}$, and shows the corresponding responses across various angular harmonics. We find that each angular harmonic contains two frequency constituents $\omega$ and $2\omega$. The $2\omega$ component arises via quadratic laminar couplings: $2\omega=\omega+\omega$, whose amplitude scales as $A_{i}^{2}$. As shown in the top panel of the figure, the shape of its angular spectrum remains unchanged over time. By contrast, the $\omega$ mode in $\ell=2$ and $\ell=4$ emerges and grows. Its angular spectrum evolves continuously (bottom panel), eventually developing greater power at larger angular scales (smaller $\ell$), indicating an inverse energy cascade in the spatial direction. Notably, this emergent growth of $\omega$ occurs regardless of the driver strength, owing to the spatial structure of the $\ell>2$ driver, which enables energy to transfer toward the largest gravitational-wave angular scale, $\ell=2$.
@@
Figure 2 provides a more detailed view of the observed behaviors for injections with $\ell=2$ (a) and $\ell=6$ (b), shown at different driving strengths $A_{i}$. In Fig. 2a, laminar modes are excited at early times, including the fourth-order mode $2\omega(=\omega+\omega+\omega-\omega)$ in $\ell=6$ (top panel), as well as the third-order mode $\omega(=\omega+\omega-\omega)$ in $\ell=6,4$ (middle and bottom panels). These modes reach their peak amplitudes just as the injected gravitational-wave driver stabilizes, shown as the gray regions. The peak amplitudes are confirmed to scale as $A_{i}^{4}$ and $A_{i}^{3}$, respectively. For small driving strengths $(A_{i}\lesssim 2\times 10^{-4})$, this early structure persists and extends to later times, suggesting a “laminar” regime in which laminar modes are present but do not interact. As the driver strength increases, a clear transition is observed: energy begins to cascade across modes, in terms of secondary emergent excitations. When the strength is within an intermediate regime, the modes grow exponentially (gray dashed lines) until saturation (open circles) and then decay. The growth becomes even faster for stronger amplitudes. Across various responses in $\ell=6$ at $2\omega$, we find that the saturation amplitude scales as $A_{i}^{4}$, while the exponential growth rate scales as $A_{i}^{2}$, implying that the nonlinear emergence involves interactions at multiple orders.
@@
Figure 2b shows how energy inversely cascades into lower $\ell$ modes while injecting $\ell=6$ gravitational waves, as discussed in Fig. 1b. We observe that the $\omega$ component in $\ell=2$ and 4 increases linearly over time, with the dashed line shown for reference. The inset plot further illustrates this linear growth. A systematic study shows that the linear growth rate scales as $A_{i}^{3}$, implying that the emergence is driven by third-order couplings.
@@
Notably, we find a similar linear growth in flat spacetime with $\ell=6$ gravitational-wave drivers. However, its growth rate is three to four orders of magnitude smaller than that around the black hole. This suggests that the inverse cascade observed in Figs. 1b and 2b is not unique to black holes but instead reflects a more generic feature of nonlinear dynamics in General Relativity, although the presence of black holes can significantly amplify the phenomenon. In comparison, the emergent behavior shown in Figs. 1a and 2a does not manifest in the flat spacetime scenario.
@@
The results in Figs. 1 and 2 suggest that the temporal and spatial inverse cascades are driven by two types of nonlinear emergence. They occur after the excitation of laminar modes (gray regions in Figs. 1a and 2a) and when the external gravitational-wave driver has stabilized to a constant amplitude. This sequence indicates that the novel excitations emerge from nonlinear interactions among these early excited laminar modes, reflecting underlying nonlinear instabilities in the spacetime. Below, we explore their possible physical origins.
@@
As mentioned, the equations of motion in General Relativity, at their core, take the schematic form $\mathcal{O}(\psi)=S(\psi)$, with $\mathcal{O}$ denoting a wave operator and $S$ a highly involved nonlinear set of terms describing gravitational interactions. The instability (emergent excitation) unearthed in Figs. 1a and 2a originates from four-mode couplings between gravitational waves, which can be expressed as
@@
where the trilinear source term $S^{(3)}$ depends on three parent modes (and possibly their complex conjugates). Focusing on the case that $\psi_{3}=\psi$ and $\omega_{1}+\omega_{2}=0$, the source term may serve as an effective anti-damping term in the wave equation, yielding $\dot{A}\propto A(A_{1}A_{2})$, or equivalently $A\propto e^{{\rm const}(A_{1}A_{2})t}$, where $A$ is the amplitude of the daughter mode $\psi$. For example, choosing the $\ell=2$ gravitational-wave driver, we have $A_{1,2}=A_{i}$ and $\omega_{1,2}=\pm\omega$; the resulting daughter mode at $2\omega$ in $\ell=6$ then grows exponentially with a rate $\propto A^{2}_{i}$ before saturation. Figure 2c shows a Feynman-type diagram for this four-mode coupling.
By contrast, the second type of instability revealed in Figs. 1b and 2b can be explained by a three-mode coupling:
@@
where $S^{(2)}$ is a bilinear second-order source that includes contributions from conjugate frequencies. Considering two parent modes $\psi_{1,2}\propto A_{1,2}e^{i\omega_{1,2}t}$, the daughter mode $\psi$ at $\omega_{1}+\omega_{2}$ is resonantly excited as $\psi\propto A_{1}A_{2}te^{i(\omega_{1}+\omega_{2})t}$. In the present case with an $\ell=6$ driver, we have $\omega_{1,2}=2\omega,-\omega$ and $A_{1,2}=A_{i}^{2},A_{i}$, yielding $A\propto A^{3}_{i}$. This “resonance”-type instability is analogous to a driven harmonic oscillator under external forcing at the same oscillation frequency. Figure 2d illustrates the corresponding Feynman-type diagram.

### v2

@@ Abstract
Gravity in nonlinear and dynamical regimes underpins spectacular astrophysical phenomena and observable consequences, from the early universe to black hole collisions. In these extreme environments, inverse energy cascades — mediated by nonlinear interactions — may help explain the near scale-invariance of cosmic structure and the simplicity of gravitational waves from binary black hole mergers. Yet the presence, characteristics, and generality of such interactions in full General Relativity remain largely unexplored. Here we show that two types of nonlinear interactions — a four-mode and a three-mode interaction — emerge in the fully nonlinear regime, and can indeed channel inverse energy cascades by inducing resonant and anti-damping (transient) nonlinear instabilities. We further demonstrate a “laminar” to “turbulent” transition for the largest-possible angular structure in General Relativity, whereas finer structures remain persistently turbulent. Our results reveal the impact and generality of these nonlinear interactions (instabilities), which can be key to understanding observations ranging from cosmological to kilometer scales. We anticipate that our work will shed new light on nonlinear gravitational phenomena and their consequences, such as constructing gravitational wave templates and testing General Relativity in the most extreme regime. Moreover, our work is a starting point for addressing nonlinear gravitational interactions using ideas and methods inspired by fluid dynamics.
@@ introduction
Understanding truly nonlinear phenomena presents significant challenges. One prominent example is furnished by fluid dynamics, whose intrinsic complexity demands ingenious strategies to unravel its rich phenomenology. A particularly powerful approach is to study systems continuously driven by an external source, thereby establishing a steady-state regime amenable to systematic analysis. Insights obtained in this regime often shed light on undriven or transient behavior. One example is the identification of direct and inverse energy cascades in 3+1 and 2+1 dimensions, respectively [24, 26, 25], which arise through mode couplings that both generate higher-order excitations and mediate energy flow across structures of different sizes, toward smaller (larger) scales with higher (lower) angular indices $\ell$ and frequencies. This understanding has provided new avenues to both control and capture the complex dynamics exhibited by turbulent regimes, e.g. [9, 31, 11, 16].
@@
Gravity, as described by General Relativity (GR), remains only partially understood in the fully nonlinear regime. While major results have been obtained for stationary spacetimes, asymptotic structures, and perturbations around special solutions, the nonlinear and dynamical aspects still pose significant challenges at the level of first principles — even though numerical simulations can now routinely model the collisions of compact binary systems and predict the resulting gravitational-wave (GW) signatures. This raises fundamental questions about the seemingly simple structure of these GWs, and how the system transitions from nonlinear to linear behavior during the post-merger stage of binary black hole (BBH) coalescence — a linear regime crucial for BH spectroscopy [2].
@@
Our main interest here is to unravel this nonlinear regime, particularly in 3+1 dimensions. Hints of exciting phenomena have been obtained through perturbative analyses in convenient settings where linear perturbations persist long enough to channel interesting nonlinear behavior. Notably, [44, 19] uncovered a parametric instability that preferentially transfers energy to larger scales. These works, in turn, were motivated by the “fluid-gravity duality”, in which certain gravitational systems in d+1 dimensions (within asymptotically anti-de Sitter spacetimes) are dual to (d-1)+1 relativistic hydrodynamics [3]. In this framework, long-wavelength perturbations of BHs correspond, on the hydrodynamic side, to scenarios where turbulence can take place [42, 8]. Explicit illustrations of such behavior have been produced through numerical simulations on both sides of the duality [17, 1]. However, such duality requires the unphysical assumption of gravity with a negative cosmological constant.
Taking inspiration from fluid dynamics, here we explore nonlinear gravity in asymptotically flat spacetimes by introducing a steady driving source that gravitationally stirs the system through incoming GWs, and capturing the full nonlinear behavior through numerical simulations of Einstein’s equations. Our approach establishes an analogous quasi–steady regime in gravity, providing a controlled setting to reveal nonlinear gravitational phenomena that would otherwise manifest transiently in non-driven scenarios. In doing so, it opens a new avenue for exploring nonlinear dynamics in GR — one that complements traditional perturbative analyses by engaging nonlinear effects directly and borrowing conceptual guidance from fluid dynamics.
Our driving waves have prescribable amplitude, frequency, and spatial extent, allowing systematic exploration of the spacetime response. By varying these parameters, we quantitatively characterize energy transfer across frequencies and spatial scales, identify dominant interaction channels and their instabilities, and distinguish linear (laminar) from nonlinear (turbulent) behavior. This framework provides a unified picture of how fully nonlinear GR drives inverse cascades and channels energy toward longer wavelengths, laying the groundwork for future analytical and phenomenological studies.
@@
We “stir” spacetimes by injecting (quasi-)steady GWs through the boundary of a finite simulation domain. This is achieved by evolving Einstein’s equations in the Generalized Harmonic Formulation, discretized using pseudo-spectral methods with constraint-preserving boundary conditions. Further details are provided in the Supplementary Material (SM).
We study scenarios that initially correspond either to flat spacetime or a non-spinning, asymptotically flat BH [^1] Linear perturbations in this regime decay faster than in the spinning case, making it more difficult to excite nonlinear behavior., where the injected waves act as a sustained source of dynamical self-interaction and turbulent gravitational phenomena. The driving is specified through a curvature scalar at the boundary [23, 27, 28, 29]:
@@
The BH is located at the center of the domain with an initial mass $M_{i}$. We adopt
@@
where the injection amplitude is linearly ramped up to $A_{i}$ over $t_{0}=600M_{i}$. This gradual turn-on is essential for stably simulating large $A_{i}$’s. The outer boundary is at $R_{\rm Bdry}=100M_{i}$. We have verified that its location does not affect conclusions.
The frequency $\omega$ is fixed and real. Wave absorption by the BH depends on the transmissivity of the potential barrier, which in linear theory rises sharply from near zero to unity as $\omega$ crosses the real part of the quasinormal-mode frequency (see SM). We find this trend persists even in our full nonlinear regime: low-$\omega$ waves are reflected, while for large $\omega$, accretion occurs, and the BH mass increases at a constant rate scaling as $A_{i}^{2}$. In the most nonlinear case, the BH mass increases by more than a factor of six. For intermediate frequencies near the quasinormal mode, the accretion rate gradually rises to a steady asymptotic value as the BH grows. Throughout, the BH spin remains consistent with zero, since the injected $(\ell=2,m=0)$ GWs carry no angular momentum.
The spacetime geometry is encoded in the metric. Its response to driving can be monitored via the “tendicity” scalar $\mathcal{E}$, which quantifies the tidal response experienced by orthogonal observers on a given worldtube [34, 32, 47, 46, 33]. This scalar is constructed from the Weyl tensor as:
@@
where $n^{a},s^{a}$ are unit normal vectors orthonormal to the spacetime foliation and the worldtube. For initially flat spacetime, we adopt a worldtube at fixed coordinate radius; while in the BH case, we choose the apparent horizon. We examine how tendicity’s harmonic components respond to the injected GW driver.
Figure 1: Stirring a non-spinning BH with a $\ell=2$ GW driver at fixed $\omega=0.5/M_{i}$ and $A_{i}=1.2\times 10^{-3}$. The $\ell=6$ response of $\mathcal{E}$ exhibits three frequency components ($3\omega,2\omega,\omega$). Their amplitude envelopes are shown in the top three panels. The $3\omega$ and $\omega$ modes first arise through third-order couplings (gray), followed by nonlinear growth and saturation of $2\omega$ and $\omega$ (orange). Bottom: time–frequency representation exhibiting an inverse cascade: $3\omega\to 2\omega\to\omega$ (yellow dashed).
Figure 2: Stirring a non-spinning BH with a $\ell=6$ GW driver at $\omega=0.5/M_{i}$ and $A_{i}=4\times 10^{-4}$. Response across angular harmonics is measured continuously in time (in blue), with representative time slices shown (dot-dashed). Each harmonic contains $2\omega$ and $\omega$. The $2\omega$ component (top) arises from quadratic couplings and maintains a stable angular spectrum, whereas the $\omega$ mode (bottom) continuously redistributes its spectrum, yielding an inverse cascade toward lower $\ell$.
Figure 3: Emergent instabilities in a stirred BH induced by $\ell=2$ GWs at $\omega=0.5/M_{i}$ with various amplitudes (colors). Shown are amplitude envelopes of $\ell=6$ ($2\omega$ and $\omega$) and $\ell=4$ ($\omega$). Modes first undergo laminar excitations via fourth- and third-order couplings (gray). At low amplitudes, the structure persists; at higher amplitudes, new excitations emerge. For intermediate amplitudes, modes grow exponentially (gray dashed) before saturation (open circles). For $2\omega$ in $\ell=6$, the exponential growth rate scales as $A_{i}^{2}$ and saturation amplitude as $A_{i}^{4}$.
Figure 4: Emergent instabilities in a stirred BH induced by $\ell=6$ GWs at $\omega=0.5/M_{i}$ with various amplitudes (colors). The $\omega$ mode in $\ell=2$ and 44 always grows linearly (dashed lines) regardless of $A_{i}$. The growth rate scales as $A_{i}^{3}$.
@@ Results
Our main conclusion is the emergence of temporal and spatial inverse energy cascades in nonlinear dynamical spacetimes. For a Schwarzschild BH, Fig. 1 illustrates this effect while injecting an $\ell=2$ GW, with $\omega=0.5/M_{i}$ and $A_{i}=1.2\times 10^{-3}$. The envelope of the response in the $\ell=6$ harmonic of $\mathcal{E}$ is shown, decomposed into its constituent frequency components. This harmonic first undergoes two nonlinear excitations at $\omega$ and $3\omega$ (gray region), arising from third-order couplings: $\omega=\omega+\omega-\omega$ and $3\omega=\omega+\omega+\omega$. Their peak amplitudes have been verified to scale as $A_{i}^{3}$. In the weakly driven regime these components evolve laminarily: their response is instantaneous, remains phase-locked to the driving wave, and does not exchange energy with other modes. These features persist over time, and the same modes are excited regardless of the injection amplitude $A_{i}$.
@@
Remarkably, for sufficiently strong injections, as in Fig. 1, novel nonlinear excitations emerge at $2\omega$ and $\omega$ sequentially in time (orange regions). These two excitations emerge, grow, and saturate after the laminar modes pass their peaks and the injected GW driver reaches its constant amplitude. In particular, the $\omega$ component first undergoes a laminar excitation and then a secondary nonlinear emergence, resulting in a two-peaked structure in its temporal evolution. The bottom panel of Fig. 1 further illustrates this phenomenology and provides more insights: the yellow dashed line, which connects the dominant laminar mode $3\omega$ to the subsequently emergent nonlinear modes $2\omega$ and $\omega$, indicates an inverse energy cascade towards smaller frequencies [^3] We focus on $\omega M_{i}=0.5$. Similar nonlinear excitations also emerge at $\omega M_{i}=0.4$ and 0.6, but not at higher frequencies e.g., $\omega M_{i}=1.5$. Further details are provided in SM. .
@@
Figure 2 considers an $\ell=6$ driver with $\omega=0.5/M_{i}$ and $A_{i}=4\times 10^{-4}$, and shows the corresponding responses across various angular harmonics. Each angular harmonic contains two frequency constituents $\omega$ and $2\omega$. The $2\omega(=\omega+\omega)$ component arises via quadratic laminar couplings, whose amplitude scales as $A_{i}^{2}$. Its angular spectrum remains unchanged over time (top). By contrast, the $\omega$ mode in $\ell=2$ and $\ell=4$ emerges and grows. The angular spectrum evolves continuously (bottom), eventually developing greater power at larger angular scales (smaller $\ell$), indicating a spatial inverse cascade. Notably, this growth of $\omega$ occurs regardless of $A_{i}$, owing to the spatial structure of the $\ell>2$ driver, which channels energy toward the largest GW angular scale, $\ell=2$.
@@
Figure 3 provides a closer look at the $\ell=2$ injections, comparing across different driving strengths $A_{i}$. Laminar modes are excited at early times, including the fourth-order mode $2\omega(=\omega+\omega+\omega-\omega)$ in $\ell=6$ (top panel), as well as the third-order mode $\omega(=\omega+\omega-\omega)$ in $\ell=6,4$ (middle and bottom panels). These modes reach peak amplitudes just as the driver stabilizes (gray regions). The peak amplitudes are confirmed to scale as $A_{i}^{4}$ and $A_{i}^{3}$, respectively. For weak driving $(A_{i}\lesssim 2\times 10^{-4})$, this early structure persists into later times, suggesting a “laminar” regime where modes are present but do not interact. As $A_{i}$ increases, a clear transition occurs: energy begins to cascade across modes, through secondary emergent excitations. For intermediate strength, the modes grow exponentially (gray dashed lines) until saturation (open circles) and then decay. The growth becomes even faster for stronger amplitudes. Across various responses in $\ell=6$ at $2\omega$, the saturation amplitude scales as $A_{i}^{4}$, while the exponential growth rate scales as $A_{i}^{2}$, implying that multiple-order interactions are involved.
@@
Figure 4 shows how energy inversely cascades into lower $\ell$ modes while injecting $\ell=6$ GWs (Fig. 2). The $\omega$ component in $\ell=2$ and 4 increases linearly over time, with the dashed line shown for reference. The inset plot further illustrates this linear growth. A systematic study finds the growth rate scales as $A_{i}^{3}$, implying third-order couplings.
@@
Notably, we find a similar linear growth in flat spacetime with $\ell=6$ drivers. However, its growth rate is three to four orders of magnitude smaller than around the BH. This suggests that the inverse cascade observed in Figs. 2 and 4 is not unique to BHs but instead reflects a more generic feature of nonlinear dynamics in GR, although the presence of BHs can significantly amplify the phenomenon. In comparison, the emergent behavior shown in Figs. 1 and 3 does not manifest in the flat spacetime scenario.
@@
Our results suggest that the temporal and spatial inverse cascades are driven by two types of nonlinear emergence. They occur after the excitation of laminar modes (gray regions in Figs. 1 and 3) and when the external GW driver has stabilized to a constant amplitude. This sequence indicates that the novel excitations emerge from nonlinear interactions among these early excited laminar modes, reflecting underlying nonlinear instabilities. Below, we explore their possible physical origins.
@@
In GR, the equations of motion schematically read $\mathcal{O}(\psi)=S(\psi)$, with $\mathcal{O}$ denoting a wave operator and $S$ a nonlinear source term. The instability unearthed in Figs. 1 and 3 originates from four-mode couplings, which can be expressed as
@@
where the trilinear source term $S^{(3)}$ depends on three parent modes (and complex conjugates). Focusing on the case that $\psi_{3}=\psi$ and $\omega_{1}+\omega_{2}=0$, the source may serve as an effective anti-damping term in the wave equation, yielding $\dot{A}\propto A(A_{1}A_{2})$, or equivalently $A\propto e^{{\rm const}(A_{1}A_{2})t}$, where $A$ is the amplitude of the daughter mode $\psi$. For example, choosing the $\ell=2$ GW driver, we have $A_{1,2}=A_{i}$ and $\omega_{1,2}=\pm\omega$; the resulting daughter mode at $2\omega$ in $\ell=6$ then grows exponentially with a rate $\propto A^{2}_{i}$ before saturation. Figure 5 (top) shows a Feynman-type diagram for this four-mode coupling.
Figure 5: Feynman-type diagram illustrating the four-mode interaction (top) underlying the emergent excitations in Fig. 3; and the three-mode interaction (bottom) in Fig. 4.
The second type of instability revealed in Figs. 2 and 4 can be explained by a three-mode coupling:
@@
where $S^{(2)}$ is a bilinear second-order source. Considering two parent modes $\psi_{1,2}\propto A_{1,2}e^{i\omega_{1,2}t}$, the daughter mode $\psi$ at $\omega_{1}+\omega_{2}$ is resonantly excited as $\psi\propto A_{1}A_{2}te^{i(\omega_{1}+\omega_{2})t}$. With an $\ell=6$ driver, we have $\omega_{1,2}=2\omega,-\omega$ and $A_{1,2}=A_{i}^{2},A_{i}$, yielding $A\propto A^{3}_{i}$. This “resonance”-type instability is analogous to a harmonic oscillator driven at its eigenfrequency. Figure 5 (bottom) illustrates the corresponding Feynman-type diagram.